OpenStudy (stamp):

[COMBINATIONS—VERIFICATION] A soccer player can choose from 4 different shirts, 2 different pairs of shorts, and 3 different pairs of socks. How many different uniform combinations can be chosen of one shirt, one pair of shorts, and one pair of socks?

OpenStudy (anonymous):

4*2*3=24

OpenStudy (anonymous):

independent events

OpenStudy (stamp):

How would you explain your reasoning for your solution

OpenStudy (anonymous):

if events are mutually independent, their probabilities multiply. also, logically, each shirt goes with a unique pair of shoe and sock; each shoe goes with a sock.

OpenStudy (anonymous):

If you really wanted to get technical: \[ \binom{4}{1}\cdot \binom{2}{1}\cdot \binom{3}{1} \]

OpenStudy (anonymous):

But since \(\binom{n}{1}=n\) it doesn't matter.

OpenStudy (anonymous):

4 x 2x3=24